Can you please help me on the following question:
A bathroom floor is being tiled. One size of square tile will cover the floor with 200 tiles. another square tile, with sides 2cm less, will use 250 tiles to cover same area. Find the size of the larger tiles to the nearest cm.
Thank you appreciate it.

Answer by stanbon(75887)

Can you please help me on the following question:
A bathroom floor is being tiled. One size of square tile will cover the floor with 200 tiles. another square tile, with sides 2cm less, will use 250 tiles to cover same area. Find the size of the larger tiles to the nearest cm.
-------
Let "s" be the length of side of a larger tile.
--------------------------------
Total Area = 200*s^2
Total Area = 250(s-2)^2
=====
200s^2 = 250(s-2)^2
-----
200s^2 = 250s^2 - 250(4s) + 250(4)
-------
50s^2 - 1000s + 1000 = 0
----
s^2 - 20s + 20 = 0
---
s = 18.94 cm
Area of larger tile:: 18.94^2 = 358.72 cm^2
-----
Cheers,
Stan H.
-------------